OpenStudy (anonymous):

Evaluate the limit as x approaches 1 of the quotient of the square root of the quantity x squared plus 3 minus 2 and the quantity x minus

OpenStudy (freckles):

\[\lim_{x \rightarrow 1} \frac{\sqrt{x^2+3}-2}{x}\]?

OpenStudy (anonymous):

x-1 in the denominator, but yes that is right

OpenStudy (freckles):

\[\lim_{x \rightarrow 1} \frac{\sqrt{x^2+3}-2}{x-1}\]

OpenStudy (freckles):

ok well pluggin in 1 on top gives us 0 and pluggin in 1 on bottom gives us 0 since we have 0/0 we have more work

OpenStudy (freckles):

I would definitely try rationalizing the numerator here

OpenStudy (anonymous):

would you simply just square everything on top

OpenStudy (freckles):

you can't do that that is changing the value of a fraction for example: \[\frac{3}{1} \neq \frac{3^2}{1^2}\] to rationalize something you might want to think about the conjugate of the thing you want to rationalize

OpenStudy (freckles):

That is multiply the top and bottom by top's conjugate

OpenStudy (freckles):

\[\frac{\sqrt{x^2+3}-2}{x-1} \cdot \frac{\sqrt{x^2+3}+2}{\sqrt{x^2+3}+2}=?\] just multiply the top out for now

OpenStudy (freckles):

Finding the limit algebraically sometimes calls for one or some of the following (and this is list is not a list of everything): A) rationalizing B) factoring C) combining fractions D) substitution E) Sometimes I referring to other limits helps also.

OpenStudy (anonymous):

okay so would the top cancel out and you be left with 1/(x-)((square root of x^2 +3)+2)

OpenStudy (freckles):

(a-b)(a+b)=a^2-b^2 <--- so it shouldn't give you 1

OpenStudy (anonymous):

Im so confused

OpenStudy (freckles):

\[(\sqrt{x^2+3}-2)(\sqrt{x^2+3}+2)=(\sqrt{x^2+3})^2-2^2=?\]

OpenStudy (anonymous):

x^2+9-4

OpenStudy (anonymous):

which eqauls x^2+5 do you set it equal to zero then and solve for x

OpenStudy (freckles):

\[(\sqrt{x^2+3}-2)(\sqrt{x^2+3}+2)=(\sqrt{x^2+3})^2-2^2=x^2+3-4\]

OpenStudy (stamp):

More simply put, recall from algebra that\[(a+b)(a-b)=a^2-b^2\]\[a=\sqrt{x^2+3} \]\[b=2\]

OpenStudy (stamp):

So\[x^2+3-4\]is\[x^2-1\]

OpenStudy (freckles):

\[\frac{\sqrt{x^2+3}-2}{x-1} \cdot \frac{\sqrt{x^2+3}+2}{\sqrt{x^2+3}+2}=\frac{x^2+3-4}{(x-1)(\sqrt{x^2+3}+2)}\]

OpenStudy (freckles):

now you should be able to see the top is factorable

OpenStudy (freckles):

and then that the top and bottom have a common factor

OpenStudy (anonymous):

do i have to multiply the two functions together in the denominater

OpenStudy (freckles):

no

OpenStudy (freckles):

just factor the top

OpenStudy (freckles):

and cancel a common factor on top and bottom

OpenStudy (stamp):

\[\frac{x^2-1}{(x-1)(\sqrt{x^2+3}+2}=\frac{(x+1)(x-1)}{(x-1)(\sqrt{x^2+3}+2}\]

OpenStudy (anonymous):

would my x-1 cancel then

OpenStudy (freckles):

yes

OpenStudy (anonymous):

then would i use one as x to find the limit

OpenStudy (freckles):

\[\frac{x+1}{\sqrt{x^2+3}+2}\] now you can plug in 1

OpenStudy (stamp):

\[\frac{x-1}{x-1}\frac{(x+1)}{\sqrt{x^2+3}+2}=\frac{x+1}{\sqrt{x^2+3}+2}\]

OpenStudy (anonymous):

then that will give me my limit?

OpenStudy (freckles):

yep